# Adaptive Alpha Weighting with PPO: Enhancing Prompt-Based LLM-Generated Alphas in Quant Trading

### v1

@@ 4.3 Comparison of PPO-Adjusted Strategy and Market Benchmark
Table 7 presents the performance metrics of both the PPO-adjusted strategy and the market benchmark. For each stock, the framework is executed ten times using non-deterministic inference to allow stochastic behavior, which is an inherent feature of the PPO adjustment and important for capturing the variability in trading outcomes. The results are reported as the mean and standard deviation in the table. Reporting both statistics provides a measure of the average performance as well as the variability across runs, thus reflecting the stability and robustness of the framework.
The PPO-adjusted strategy shows strong performance across most companies compared to their respective market benchmarks. Apple, HSBC, Pepsi, and Tencent all achieved significantly higher cumulative returns, suggesting that the strategy was effective in capturing upside potential. Toyota, however, experienced a relatively lower return, underperforming both the other companies and its benchmark. This may indicate that the strategy struggled in the Japanese market or that market conditions during the test period were less favorable.
When looking at risk-adjusted performance through the Sharpe ratio, the PPO-adjusted strategy maintained consistently high values for Apple, Pepsi, and Tencent. This implies not only strong returns but also relatively low volatility in these cases. On the other hand, Toyota had the lowest Sharpe ratio, reinforcing the poor performance observed in its cumulative return. Drawdown values were low for most companies, indicating limited downside risk. However, Toyota again stood out with a larger drawdown, which aligns with its lower return and Sharpe ratio. Overall, the PPO strategy appears to be effective in various markets, though its success may vary depending on the specific company or market environment.
The standard deviations provide insight into the consistency of the results. For Apple and Pepsi, the standard deviations are very small relative to the mean, indicating highly consistent performance. HSBC and Tencent exhibit moderate variability, with fluctuations larger than those of the U.S. stocks. In contrast, Toyota shows relatively large standard deviations compared to its mean values, suggesting that its results are less stable. This instability may be due to several factors: the LLM-generated alphas may capture weaker or noisier signals for Toyota, making the PPO-adjusted weights fluctuate more; Toyota trades in the Nikkei 225, which can have higher volatility and lower liquidity compared to the U.S. markets.
4.4 Comparison of PPO-Adjusted Strategy and Equal-Weighted Strategy
Table 8 shows the results of an equal-weighted strategy constructed using the same set of alphas. In general, the PPO-adjusted strategy demonstrates superior performance across most stocks except Toyota when compared to the equal-weighted strategy. Specifically, for most of the time, the PPO-adjusted approach achieves higher cumulative returns and Sharpe ratios, indicating both stronger profitability and more efficient risk-adjusted performance. Furthermore, the PPO-adjusted strategy exhibits lower maximum drawdowns, reflecting improved downside protection.
In contrast, the equal-weighted strategy produces mixed results, with lower returns and higher drawdowns for most stocks, highlighting the limitations of assigning equal weight to all alphas without considering their predictive contribution and market dynamics. These comparisons suggest that the PPO-adjusted framework effectively identifies and emphasizes more informative alphas, leading to a more robust and stable trading performance relative to the equal-weighted approach.
Table 7: PPO-Adjusted Strategy vs. Market Benchmark (with 50 Alphas), Mean (Standard Deviation)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return 1.6817 (0.0619) 0.4657 (0.0554) 0.6272 (0.0331) 0.6245 (0.0632) 0.0299 (0.0398)
Sharpe Ratio 1.9998 (0.0169) 0.8131 (0.0691) 1.4319 (0.0434) 1.1440 (0.1052) 0.1011 (0.0786)
Drawdown -0.0101 (0.0007) -0.2259 (0.0099) -0.0067 (0.0001) -0.0810 (0.0208) -0.2801 (0.0076)
Market Return 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.4081 (Nikkei 225)
Table 8: Equal-Weighted Strategy (with 50 Alphas)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return -0.3200 -0.9059 -0.1726 -0.7074 0.1026
Sharpe Ratio -0.3601 -0.3580 -0.2879 -0.2032 0.1586
Drawdown -0.3663 -0.9489 -0.2922 -0.8741 -0.3571
Market Return 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.4081 (Nikkei 225)
Table 9: PPO-Adjusted Strategy Using Low-Correlation Alphas vs. Market Benchmark, Mean (Standard Deviation)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return 1.6641 (0.1755) 0.4983 (0.0660) 0.6319 (0.0775) 0.5956 (0.0722) 0.0842 (0.0579)
Sharpe Ratio 2.0002 (0.0692) 0.8388 (0.0776) 1.4259 (0.0622) 1.1191 (0.1018) 0.2072 (0.1083)
Drawdown -0.0113 (0.0016) -0.2210 (0.0121) -0.0066 (0.0002) -0.0810 (0.0183) -0.2344 (0.0288)
Market Return 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.4081 (Nikkei 225)
Table 10: PPO-Adjusted Strategy Using High-Contribution Alphas vs. Market Benchmark, Mean (Standard Deviation)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return 1.6072 (0.1093) 0.5199 (0.0991) 0.6023 (0.0441) 0.5634 (0.1200) 0.0741 (0.0395)
Sharpe Ratio 1.9890 (0.0626) 0.8791 (0.1215) 1.3996 (0.0296) 1.0620 (0.1735) 0.1894 (0.0765)
Drawdown -0.0113 (0.0016) -0.2286 (0.0253) -0.0076 (0.0022) -0.0919 (0.0263) -0.2625 (0.0217)
Market Return 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.4081 (Nikkei 225)
Table 11: PPO-Adjusted Strategy Using 30 Randomly Selected Alphas vs. Market Benchmark, Mean (Standard Deviation)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return 1.7276 (0.0647) 0.4853 (0.0451) 0.6176 (0.0225) 0.6041 (0.0755) 0.0410 (0.0291)
Sharpe Ratio 2.0133 (0.0255) 0.8399 (0.0534) 1.4271 (0.0367) 1.1129 (0.1132) 0.1238 (0.0565)
[…]

### v2

@@ 4.4 Comparison of PPO-Adjusted Strategy and Benchmarks
Table 7 presents the performance metrics of both the PPO-adjusted strategy and different benchmarks. For each stock, the PPO framework is executed ten times using non-deterministic inference to allow stochastic behavior, which is an inherent feature of the PPO adjustment and important for capturing the variability in trading outcomes. The results are reported as the mean and standard deviation in the table. Reporting both statistics provides a measure of the average performance as well as the variability across runs, thus reflecting the stability and robustness of the framework.
Overall, the PPO strategy delivers more favorable risk-adjusted outcomes across most stocks in terms of Sharpe Ratio. Although the Buy-and-Hold (B&H) strategy achieves higher cumulative returns in most cases, such as Toyota, Airbus, and Netflix, these gains are accompanied by higher risk exposure. In contrast, the PPO strategy records consistently higher Sharpe Ratios for nearly all stocks, indicating a more stable return profile.
A notable feature of the PPO strategy is its strong control of downside risk. For all ten stocks, the maximum drawdown remains very small, often below $1\%$. This behavior suggests that the agent has learned to reduce exposure during unfavorable market conditions, including periods of elevated volatility or persistent price declines. By comparison, the Equal-Weighted (EW) and Momentum (MOM) strategies experience substantial drawdowns, exceeding $50\%$ in several cases, such as Apple and Tencent. These results highlight the PPO agent’s emphasis on capital preservation rather than aggressive return seeking.
Results of significance tests
To examine whether the observed performance differences are statistically meaningful, Table 8 presents the Diebold–Mariano (DM) test results based on daily returns. In most comparisons, the DM statistics are negative, indicating lower forecast errors for the PPO strategy relative to the benchmark methods.
When compared with the weaker baselines, namely the EW and MOM strategies, the PPO strategy shows statistically significant improvements at the $1\%$ or $5\%$ level for most of the stocks. This indicates that the learned policy captures market information that simple heuristic rules fail to reflect.
In contrast, the DM test results against the B&H strategy are less frequently significant. For example, the $p$-values for Apple ($p=0.071$) and HSBC ($p=0.224$) do not reach conventional significance levels. This outcome suggests that during extended upward market phases, the daily return patterns of the PPO strategy and B&H can be similar, even though their risk characteristics differ. However, according to the Sharpe ratio bootstrap test result in Table 9, the PPO strategy yields higher Sharpe ratios than Buy-and-Hold for most stocks. Statistically significant improvements are observed in seven out of ten cases, while the remaining stocks show no significant differences. These findings indicate that PPO offers consistent risk-adjusted performance gains, although the magnitude and significance of the improvement depend on the underlying asset.
A possible explanation for the lack of statistically significant DM test results when comparing the PPO strategy with the Buy-and-Hold (B&H) benchmark relates to differences in trading activity. Figure 5 illustrates the portfolio value, position exposure, and alpha signal strength over the test period for Airbus. As is shown in Figure 5, the PPO agent adopts a highly selective execution policy and frequently remains market-neutral for extended periods. As a result, a substantial portion of the return series consists of zero or near-zero values corresponding to no-trade days.
Since the Diebold–Mariano test assigns equal weight to all observations, including periods without active trading, strategies with sparse execution may experience a dilution effect in the test statistic. From a statistical perspective, this selective participation effectively reduces the informative sample size relative to continuously invested benchmarks such as B&H. Consequently, the resulting $p$-values may be more conservative, potentially understating the statistical significance of the PPO agent’s predictive signal despite its superior risk-adjusted performance.
Finally, comparisons with the Random Entry/Exit Baseline (RB) show strong statistical evidence in favor of the PPO strategy for most stocks, including Toyota and Tencent (both with $p<0.001$). This confirms that the observed performance is not driven by random trading behavior but instead reflects the effect of policy learning.
The consistency between the performance indicators in Table 7 and the DM test results in Table 8 points to a stable and reliable trading policy. Rather than maximizing raw returns, the PPO agent appears to focus on improving the Sharpe Ratio by limiting downside variability. Even in cases where cumulative returns fall below those of the B&H strategy, such as Airbus, the substantially lower drawdowns lead to a more efficient risk–return trade-off. Overall, these findings suggest that reinforcement learning can be an effective tool for constructing trading strategies that balance profitability with robustness under volatile market conditions.
Figure 5: Performance Analysis and Execution Dynamics of the PPO Strategy for Pepsi Stock. The panels illustrate the cumulative portfolio value (top), the volatility-adjusted position exposure $p_{t}$ (middle), and the composite alpha signal strength relative to decision thresholds (bottom).
Table 7: Comprehensive Performance Analysis: PPO Strategy vs. EW, Buy & Hold, Momentum, and Random Entry/Exit Baselines. The best result under each metric for each stock is highlighted in bold.
Stock Cumulative Return Sharpe Ratio Max Drawdown
PPO (Std) EW B&H MOM RB (Std) PPO (Std) EW B&H MOM RB (Std) PPO (Std) EW B&H MOM RB (Std)
Apple 1.6817 (0.0619) 0.3200-0.3200 0.3082 0.4797-0.4797 0.1427 (0.1640) 1.9998 (0.0169) 0.3601-0.3601 0.8194 0.9119-0.9119 0.5283 (0.5358) $\mathbf{-0.0101}$ (0.0007) 0.3663-0.3663 0.1732-0.1732 0.5214-0.5214 0.1726-0.1726 (0.0484)
HSBC 0.5685 (0.0259) 0.9059-0.9059 0.6161 0.5681-0.5681 0.2899 (0.1671) 2.4113 (0.0558) 0.3580-0.3580 1.6087 1.2032-1.2032 1.1838 (0.5961) $\mathbf{-0.0220}$ (0.0099) 0.9489-0.9489 0.1425-0.1425 0.6479-0.6479 0.1179-0.1179 (0.0333)
Pepsi 0.6272 (0.0331) 0.1726-0.1726 0.1170 0.0231 0.0556 (0.1012) 1.4319 (0.0434) 0.2879-0.2879 0.5220 0.2131 0.3292 (0.5531) $\mathbf{-0.0067}$ (0.0001) 0.2922-0.2922 0.1694-0.1694 0.2698-0.2698 0.1305-0.1305 (0.0423)
Tencent 0.6245 (0.0632) 0.7074-0.7074 0.1752 0.4757-0.4757 0.0999 (0.2806) 1.1440 (0.1052) 0.2032-0.2032 0.4680 0.5754-0.5754 0.2860 (0.5614) $\mathbf{-0.0810}$ (0.0208) 0.8741-0.8741 0.3362-0.3362 0.6751-0.6751 0.2930-0.2930 (0.0716)
Toyota 0.0638 (0.0267) 0.1026 0.8742 0.2464-0.2464 0.3956 (0.2490) 1.2786 (0.3810) 0.1586 1.6646 0.5608-0.5608 1.1780 (0.6317) $\mathbf{-0.0223}$ (0.0062) 0.3571-0.3571 0.1420-0.1420 0.2820-0.2820 0.1422-0.1422 (0.0365)
Airbus 0.0148 (0.0079) 0.3687-0.3687 0.6646 0.1428-0.1428 0.2786 (0.1269) 1.3899 (0.5449) 0.1245-0.1245 2.4272 0.6256-0.6256 1.6558 (0.6503) $\mathbf{-0.0061}$ (0.0042) 0.5228-0.5228 0.0757-0.0757 0.2901-0.2901 0.0830-0.0830 (0.0241)
Exxon Mobil 0.0307 (0.0175) 0.1131-0.1131 0.4597 0.0684 0.1765 (0.1374) 1.5525 (0.2590) 0.0941-0.0941 1.5336 0.3741 0.9129 (0.6201) $\mathbf{-0.0030}$ (0.0027) 0.2994-0.2994 0.1567-0.1567 0.1358-0.1358 0.1455-0.1455 (0.0388)
Petrobras 0.0945 (0.0356) 0.1403-0.1403 0.4659 0.4124-0.4124 0.2553 (0.2505) 1.5357 (0.2272) 0.0736 1.1112 1.1008-1.1008 0.8617 (0.7045) $\mathbf{-0.0102}$ (0.0042) 0.5745-0.5745 0.3154-0.3154 0.5565-0.5565 0.2313-0.2313 (0.0783)
Netflix 0.3085 (0.0009) 0.3756-0.3756 0.6215 0.1572-0.1572 0.2389 (0.2541) 1.9010 (0.0050) 0.3469-0.3469 1.3044 0.2247-0.2247 0.7507 (0.6150) $\mathbf{-0.049}$ (0.0004) 0.4525-0.4525 0.2010-0.2010 0.3524-0.3524 0.1865-0.1865 (0.0484)
InfuSystem 0.0202 (0.0182) 0.6989-0.6989 0.0912 0.0169-0.0169 0.0666 (0.2297) 0.8970 (0.6531) 0.5680-0.5680 0.4121 0.2005 0.2967 (0.5985) $\mathbf{-0.0101}$ (0.0045) 0.7443-0.7443 0.3405-0.3405 0.3225-0.3225 0.2583-0.2583 (0.0751)
Table 8: Diebold-Mariano (DM) Test Results: PPO Strategy vs. Benchmarks. The PPO results are derived from the average daily returns across multiple independent training runs (seeds).
Stock PPO vs. B&H PPO vs. Random PPO vs. EW PPO vs. MOM
DM Stat $p$-value DM Stat $p$-value DM Stat $p$-value DM Stat $p$-value
Apple 1.81-1.81 0.071 3.26-3.26** 0.001 4.96-4.96** 0.000 2.49-2.49* 0.013
HSBC 1.22-1.22 0.224 2.46-2.46* 0.014 2.43-2.43* 0.015 2.86-2.86* 0.004
Pepsi 1.34-1.34 0.180 1.79-1.79 0.074 2.59-2.59** 0.009 1.51-1.51 0.131
Tencent 2.25-2.25* 0.024 3.49-3.49** 0.000 0.37-0.37 0.710 3.07-3.07** 0.002
Toyota 1.34-1.34 0.180 3.53-3.53** 0.000 1.38-1.38 0.168 3.76-3.76** 0.000
Airbus 0.54-0.54 0.589 0.20-0.20* 0.045 3.94-3.94** $0.001$ 1.74-1.74 0.084
Exxon Mobil 0.49 0.625 0.40 0.689 1.96-1.96* 0.050 0.27-0.27 0.783
Petrobras 0.14 0.888 1.78-1.78 0.075 2.73-2.73** 0.006 2.05-2.05* 0.040
Netflix 0.61-0.61 0.546 1.23-1.23 0.222 3.27-3.27** 0.002 2.16-2.16* 0.032
InfuSystem 1.18-1.18 0.240 2.06-2.06* 0.041 5.27-5.27** 0.000 0.43-0.43 0.665
Note: * and ** indicate significance at the 5% and 1% levels, respectively.
A negative DM stat indicates PPO outperforms the benchmark.
Table 9: Relative Sharpe Ratio Bootstrap Test Results (PPO vs. Buy-and-Hold). The Sharpe ratio difference is defined as $\Delta SR=SR_{\text{PPO}}-SR_{\text{B\&H}}$. PPO results are computed using average daily returns across multiple independent training runs (seeds).
Stock Sharpe Difference 95% CI $p$-value
Apple 1.7695* [0.1787, 3.4424] 0.0296
HSBC 0.7928* $[0.4321,2.2641]$ 0.0006
Pepsi 0.3365* [0.0010, 1.9168] 0.0494
Tencent 1.3282 [0.6347,3.1084][-0.6347,3.1084] 0.1820
Toyota 0.4890* [0.2533, 1.2456] 0.0010
Airbus 0.1380-0.1380 [2.9630,2.2057][-2.9630,2.2057] 0.7814
Exxon Mobil 0.1628* [0.0857, 1.1749] 0.0370
Petrobras 0.0903 [1.3394,0.6880][-1.3394,0.6880] 0.8062
Netflix 0.6016* [0.1274, 1.3747] 0.0174
InfuSystem 1.0333* [0.5141, 2.4829] <0.0001<0.0001
Note: indicates statistical significance at the 5% level.
Note: The 95% confidence intervals are obtained via block bootstrap resampling.
Statistical significance is determined by whether the confidence interval excludes zero.
A positive Sharpe difference indicates that the PPO strategy outperforms Buy-and-Hold.
PPO execution metrics
Table 10 summarizes the execution-related properties of the PPO strategy. These indicators provide useful context for understanding how the agent translates its learned policy into actual trading behavior.
Across the ten stocks, the observed win rates range from $14.16\%$ to $44.19\%$. Although these values are relatively low in absolute terms, they should be viewed together with the strong risk-adjusted performance and limited drawdowns reported in Table 7. This pattern suggests that the PPO agent does not depend on a high frequency of profitable trades. Instead, it appears to limit losses on unfavorable positions while allowing successful trades to run, thereby achieving stable performance despite a low hit ratio.
The average duration of open positions differs substantially across stocks, indicating flexible timing behavior. For more volatile or growth-oriented stocks such as Netflix, the agent tends to hold positions for longer periods, with an average duration of $14.05$ days. In contrast, for assets like Airbus and Toyota, the holding periods are much shorter, typically around $1.0$ to $1.3$ days. This difference suggests that the PPO policy adapts its trading horizon to asset-specific dynamics rather than applying a uniform temporal rule.
Furthermore, the strategy exhibits moderate trading activity. Average daily turnover remains low for several stocks, including Netflix ($0.0090$) and HSBC ($0.0348$), and reaches its highest level for Airbus ($0.2007$). Such turnover characteristics are important in practice, as they reduce the likelihood that transaction costs and market impact substantially diminish the strategy’s realized returns.
Table 10: Practical Trading Metrics Sample of the PPO-Adjusted Strategy across Ten Stocks
Stock Avg. Win Rate Avg. Position Duration (Days) Avg. Daily Turnover
Apple 0.2582 3.8407 0.0616
HSBC 0.3704 8.0426 0.0348
Pepsi 0.1416 6.0783 0.0483
Tencent 0.1993 4.6610 0.0676
Toyota 0.2911 1.3390 0.1015
Airbus 0.4270 1.0001 0.2007
Exxon Mobil 0.3783 1.0553 0.1795
Petrobras 0.2360 1.6084 0.0989
Netflix 0.4419 14.0526 0.0090
InfuSystem 0.1729 1.8601 0.0613
5 Further Discussion and Analysis
In this section, we select a subset of five stocks to conduct further analysis, such as the ablation analysis, wark-forward optimization analysis, impact of alpha selection settings, prompt information and sentiment on trading performance.
5.1 Ablation Analysis of Alpha Sources
To examine whether the use of a large language model contributes incremental value beyond conventional feature engineering, we conduct a controlled ablation analysis based on a subset of five stocks that contrasts LLM-generated alphas with a standardized set of human-crafted alphas. The latter consists of 101 manually designed factors Kakushadze (2016) that are widely used in quantitative finance and serve as a representative benchmark for traditional alpha construction.
In order to ensure a fair comparison, the two alpha sources are evaluated under an identical experimental setup. Specifically, both sets are derived from the same underlying price and volume data, and share the same normalization procedures. Following this preprocessing stage, the retained alphas are fed into the same PPO-based optimization framework. All model architectures, hyperparameters, and reward definitions are kept fixed across experiments.
Since the number of LLM-generated alphas differs from the size of the human-crafted pool, we control for feature dimensionality by selecting a matched subset of 50 human-crafted alphas. This design choice prevents differences in performance from being driven by unequal feature counts and allows the comparison to focus on the alpha generation mechanism itself. Table 11 shows an example of five human-crafted alphas used for comparison.
By isolating the source of the alphas while holding all other components constant, this ablation study directly assesses whether the LLM provides additional predictive and economic value beyond established human-designed factors.
Table 13 shows the performance comparison between human-crafted and LLM-generated alphas under the same PPO strategy. In general, the LLM-generated alphas have better performance for almost all stocks except Tencent, which suggests that LLM-generated alphas can serve as a reliable and effective alternative to manually designed signals within the PPO framework.
Table 11: Example of Human-Crafted Alphas Used for Comparison
Alpha Definition
Alpha 1 $\mathrm{Rank}\!\left(\mathrm{TsArgMax}\!\left(\left(\begin{cases}\mathrm{Std}(\mathrm{Ret},20),&\text{if }\mathrm{Ret}<0\\ \mathrm{Close},&\text{otherwise}\end{cases}\right)^{2},5\right)\right)$
Alpha 2 $\mathrm{Corr}\!\left(\mathrm{Rank}\!\left(\Delta\log(\mathrm{Vol}),2\right),\mathrm{Rank}\!\left(\frac{\mathrm{Close}-\mathrm{Open}}{\mathrm{Open}}\right),6\right)$
Alpha 3 $\mathrm{Corr}\!\left(\mathrm{Rank}(\mathrm{Open}),\mathrm{Rank}(\mathrm{Vol}),10\right)$
Alpha 4 $\mathrm{TsRank}\!\left(\mathrm{Rank}(\mathrm{Low}),9\right)$
Alpha 5 $\mathrm{Rank}\!\left(\mathrm{Open}-\frac{1}{10}\sum_{t=1}^{10}\mathrm{VWAP}\right)\cdot\left(-\,\left|\mathrm{Rank}(\mathrm{Close}-\mathrm{VWAP})\right|\right)$
Table 12: Comparison between Human-Crafted and LLM-Generated Alphas under the PPO Strategy for ten runs. Values are reported as Mean (Standard Deviation).
Stock Metric Human-Crafted LLM-Generated
Apple Cumulative Return 0.0085-0.0085 (0.0128) 1.6817 (0.0619)
Sharpe Ratio 0.5914-0.5914 (0.9719) 1.9998 (0.0169)
Max Drawdown 0.0125-0.0125 (0.0100) 0.0101-0.0101 (0.0007)
HSBC Cumulative Return 0.0040 (0.0074) 0.5685 (0.0259)
Sharpe Ratio 0.3402 (0.9468) 2.4113 (0.0558)
Max Drawdown 0.0050-0.0050 (0.0041) 0.0220-0.0220 (0.0099)
Pepsi Cumulative Return 0.0027-0.0027 (0.0313) 0.6272 (0.0331)
Sharpe Ratio 0.0793-0.0793 (0.4538) 1.4319 (0.0434)
Max Drawdown 0.0264-0.0264 (0.0095) 0.0067-0.0067 (0.0001)
Tencent Cumulative Return 0.0495 (0.0235) 0.6245 (0.0632)
Sharpe Ratio 2.0123 (0.5520) 1.1440 (0.1052)
Max Drawdown 0.0051-0.0051 (0.0032) 0.0810-0.0810 (0.0208)
Toyota Cumulative Return 0.0187-0.0187 (0.0285) 0.0638 (0.0267)
Sharpe Ratio 0.5228-0.5228 (0.6423) 1.2786 (0.3810)
Max Drawdown 0.0354-0.0354 (0.0249) 0.0223-0.0223 (0.0062)
5.2 Evaluation of different RL algorithms within the proposed framework
In this subsection, we investigate whether PPO is the only model suitable for the proposed framework. We compare PPO with A2C, SAC, and TD3 under the same environmental setup. A brief overview of these algorithms is presented below.
  • Advantage Actor–Critic (A2C) Mnih et al. (2016): A2C is a synchronous actor–critic method that updates the policy using on-policy samples. By estimating the advantage function, it reduces the variance of policy updates and provides a straightforward and stable baseline for policy-gradient-based trading strategies.
  • Soft Actor–Critic (SAC) Haarnoja et al. (2018): SAC is an off-policy algorithm that incorporates an entropy term into the objective function. This design encourages broader exploration and can lead to more stable learning, making SAC suitable for evaluating risk-aware and conservative trading behaviors under the same environment.
  • Twin Delayed Deep Deterministic Policy Gradient (TD3) Fujimoto et al. (2018): TD3 is an off-policy method that reduces value overestimation by using two critic networks and delaying policy updates. These mechanisms help stabilize learning in continuous action spaces, which is beneficial when controlling position sizes in trading tasks.
Table 13 summarizes the performance of PPO and several baseline reinforcement learning algorithms across five selected stocks. Overall, the results indicate that model performance varies across both stocks and evaluation metrics, and no single method consistently achieves the best outcomes in all cases.
In terms of cumulative return, PPO tends to obtain higher values for Apple, Pepsi, and Tencent. However, SAC, TD3, and A2C often achieve higher Sharpe ratios for the same stocks. This pattern may suggest a trade-off between return and risk. A possible explanation is that PPO adopts relatively more active trading behaviors, which can increase return variability, while the other algorithms may favor more stable policies that lead to smoother returns. As a result, higher cumulative returns do not necessarily translate into higher risk-adjusted performance.
A notable difference is observed for Pepsi, where PPO produces a substantially larger cumulative return than the other methods. This gap may be related to the characteristics of the Pepsi stock during the evaluation period. It is possible that PPO is more sensitive to short-term signals or temporary trends, allowing it to capture certain profitable opportunities. In contrast, off-policy methods such as SAC and TD3 may converge to more conservative strategies, which could limit their exposure and reduce returns in this particular case.
In terms of downside risk, the maximum drawdown values are generally comparable across methods for most stocks. PPO exhibits slightly larger drawdowns in some cases, which is consistent with its higher return volatility. However, these differences remain relatively small and do not indicate excessive risk-taking.
Furthermore, the relatively small standard deviations observed across repeated experiments may be attributed to the risk-control mechanisms embedded in the trading environment. In particular, volatility targeting and bounded position sizing constrain the magnitude of portfolio exposure, which can limit performance variability across different runs. In addition, all models are evaluated on the same historical price trajectory under identical experimental settings, with differences arising mainly from policy initialization. Under such conditions, outcome variability is expected to remain limited.
In conclusion, this analysis suggests that PPO is not strictly required and may be substituted with other reinforcement learning models using the same environmental setup.
Table 13: Comparison between PPO and other RL algorithms. Values are reported as mean with standard deviation shown below.
Stock / Metric PPO SAC TD3 A2C
Apple
Cumulative Return 1.6817 (0.0619) 0.5621 (0.0075) 0.5845 (0.0085) 0.5619 (0.0146)
Sharpe Ratio 1.9998 (0.0169) 2.6061 (0.0113) 2.6236 (0.0126) 2.5899 (0.0266)
Max Drawdown 0.0101-0.0101 (0.0007) 0.0099-0.0099 (0.0003) 0.0103-0.0103 (0.0001) 0.0101-0.0101 (0.0002)
HSBC
Cumulative Return 0.5685 (0.0259) 0.5563 (0.0382) 0.5605 (0.0039) 0.5563 (0.0145)
Sharpe Ratio 2.4113 (0.0558) 2.3718 (0.0958) 2.3812 (0.0113) 2.3746 (0.0331)
Max Drawdown 0.0220-0.0220 (0.0007) 0.0220-0.0220 (0.0009) 0.0220-0.0220 (0.0001) 0.0220-0.0220 (0.0001)
Pepsi
Cumulative Return 0.6272 (0.0331) 0.0568 (0.0003) 0.0567 (0.0002) 0.0569 (0.0002)
Sharpe Ratio 1.4319 (0.0434) 0.8597 (0.0045) 0.8587 (0.0031) 0.8617 (0.0030)
Max Drawdown 0.0067-0.0067 (0.0001) 0.0010-0.0010 (0.0001) 0.0012-0.0012 (0.0001) 0.0010-0.0010 (0.0001)
Tencent
Cumulative Return 0.6245 (0.0632) 0.3270 (0.0050) 0.3262 (0.0011) 0.3251 (0.0021)
Sharpe Ratio 1.1440 (0.1052) 1.8083 (0.0216) 1.8054 (0.0051) 1.8011 (0.0090)
Max Drawdown 0.0810-0.0810 (0.0208) 0.0548-0.0548 (0.0001) 0.0541-0.0541 (0.0003) 0.0541-0.0541 (0.0005)
Toyota
Cumulative Return 0.0638 (0.0267) 0.0722 (0.0111) 0.0579 (0.0085) 0.0821 (0.0265)
Sharpe Ratio 1.2786 (0.3810) 1.2039 (0.1694) 1.2012 (0.1665) 1.6646 (0.3856)
Max Drawdown 0.0223-0.0223 (0.0062) 0.0242-0.0242 (0.0038) 0.0211-0.0211 (0.0028) 0.0158-0.0158 (0.0032)
5.3 Walk-Forward Optimization Analysis
In the original experiments, the dataset is split into a fixed training period and a subsequent testing period. Models are trained once on historical data and evaluated on unseen future data. To further evaluate the robustness of the proposed framework under different market conditions, we additionally conduct a walk-forward optimization analysis (Figure 6). Unlike the static train–test split, walk-forward optimization repeatedly retrains the model on a rolling training window and evaluates it on the immediately following out-of-sample period.
Financial markets exhibit strong non-stationarity and regime shifts. A single fixed split may overestimate performance if the training and testing periods share similar market conditions. Walk-forward optimization provides a more realistic evaluation by mimicking practical deployment, where models are periodically retrained using the most recent data. For walk-forward optimization, performance metrics are computed for each out-of-sample window and then averaged across all windows to obtain a stable estimate.
In this paper, we adopt a walk-forward evaluation scheme with a rolling training window of two years ($252\times 2$ days), followed by a one-year (252 days) out-of-sample testing period. The window is advanced by six months (126 days) at each step, allowing partial overlap between consecutive test periods. This design aims to balance estimation stability and adaptability to evolving market conditions, while maintaining a strict separation between training and testing data. The PPO strategy is evaluated against a Buy-and-Hold strategy and a Momentum strategy within the same test window.
Table 14 reports the walk-forward out-of-sample results for PPO, Buy-and-Hold, and a simple Momentum strategy across five stocks. Overall, the three methods show clearly different performance patterns under the same rolling evaluation setting.
Similar to the fixed train-test split scenario, across all stocks, Buy-and-Hold and Momentum strategies generally achieve higher total returns than PPO. This behavior is consistent with their continuous or near-continuous market exposure during the test windows. In contrast, PPO produces lower average returns, but it is associated with substantially lower annualized volatility and smaller maximum drawdowns. As a result, PPO often exhibits competitive, and in some cases higher, Sharpe ratios compared with the benchmark strategies.
The Momentum strategy behaves similarly to Buy-and-Hold for some stocks, such as Apple and Pepsi, where momentum signals remain positive for long periods. In these cases, the Momentum strategy maintains sustained exposure, leading to comparable returns, volatility, and drawdowns. For other stocks, such as HSBC and Toyota, Momentum shows weaker performance, which may indicate that trend persistence is less stable in these markets. However, this interpretation remains tentative, as the exact behavior depends on the underlying price dynamics within each test window.
The observed return–risk trade-off between PPO and the benchmark strategies appears reasonable in the walk-forward setting. Buy-and-Hold and Momentum strategies prioritize market exposure, which can lead to higher cumulative returns but also larger drawdowns and volatility. PPO, on the other hand, adopts a more cautious trading behavior, adjusting positions over time and limiting exposure during certain periods. This difference may help explain why PPO achieves lower drawdowns and volatility, even when raw returns are smaller.
The lower win rates observed for PPO do not necessarily imply inferior performance. A possible explanation is that PPO trades less frequently or remains inactive during many periods, which reduces the proportion of positive-return days but may help control risk. Similarly, the relatively small standard deviations reported for PPO across windows may be related to its more stable position adjustments.
Overall, the results suggest that PPO and traditional benchmark strategies exhibit distinct performance characteristics under walk-forward evaluation. PPO appears to emphasize stability and risk control, while Buy-and-Hold and Momentum focus more on sustained market exposure. These differences highlight the importance of evaluating trading strategies from both return and risk perspectives rather than relying on a single metric. Furthermore, the performance of PPO is relatively consistent with the fixed train-split result shown in Table 7.
Figure 6: Timeline illustration of walk-forward optimization. Each window consists of a rolling in-sample training period followed by an out-of-sample testing period. The window advances forward in time while maintaining a fixed training and testing length.
Table 14: Walk-forward out-of-sample performance comparison between PPO and Buy-and-Hold. Values are reported as mean (standard deviation) over all rolling test windows.
Stock Method Total Return Ann. Vol. Sharpe Max DD Win Rate Daily Turnover
Apple PPO 0.0479 (0.0563) 0.0474 (0.0520) 1.5530 (0.8644) $\mathbf{-0.0040}$ (0.0056) 0.1392 (0.1181) 0.0427 (0.0341)
Buy-and-Hold 0.3573 (0.3912) 0.3160 (0.0837) 0.9581 (0.7357) 0.2502-0.2502 (0.0732) 0.5291 (0.0293) 0.0000 (0.0000)
Momentum 0.3665 (0.4059) 0.3156 (0.0835) 0.9684 (0.7410) 0.2530-0.2530 (0.0695) 0.5290 (0.0296) 1.000 (<0.0001<0.0001)
HSBC PPO 0.0379 (0.1219) 0.0913 (0.2643) 1.1868 (2.1042) $\mathbf{-0.0256}$ (0.0510) 0.1609 (0.1398) 0.0566 (0.0559)
Buy-and-Hold 0.1855-0.1855 (0.1814) 0.2359 (0.1070) 0.7394-0.7394 (0.8155) 0.7082-0.7082 (0.3647) 0.4843 (0.0419) 0.0000 (0.0000)
Momentum 0.0600-0.0600 (0.1437) 0.0639 (0.1030) 0.8284-0.8284 (0.9728) 0.0941-0.0941 (0.1660) 0.1543 (0.2322) 0.3333 (0.5000)
Pepsi PPO 0.0220 (0.0657) 0.0389 (0.0463) 0.8481 (1.5084) $\mathbf{-0.0109}$ (0.0291) 0.1352 (0.1435) 0.0403 (0.0394)
[…]